# Exploring Parent-Child Perceptions on Safety in Generative AI: Concerns, Mitigation Strategies, and Design Implications

### v1

@@ 6.1 Design Implications
There are existing AI risk management frameworks, policies, and regulations for children, such as those provided by the United Nations Children’s Fund (UNICEF), the United Nations Convention on the Rights of the Child (UNCRC), and the Age Appropriate Design Code (AADC) [13]. However, finding this compliance information for specific products is difficult for users, especially children. We recommend that GAI practitioners combine these AI risk management frameworks with users’ perceptions of risks to create an easy-to-understand child risk and protection disclosure on product pages. User feedback and priorities on GAI risks and protections can guide the design of these disclosures, influencing the sequence and format of information. Additionally, this feedback can inform policymakers in creating more effective and user-centered regulations. To make this disclosure easily understandable for both parents and children, GAI practitioners could draw inspiration from label designs used for children’s products in everyday life. Emami-Naeini et al. have prototyped security and privacy labels for IoT devices, which could serve as a model for creating similar labels for GAI platforms. This would help parents and children by providing clear, accessible information about the safety and privacy measures in place for each platform, lowering the barrier for them to make informed decisions.
@@
While existing parental control features are useful, they are not entirely suitable for managing appropriate content on Generative AI (GAI) platforms. The primary challenge stems from the nature of GAI, which generates real-time outputs rather than providing pre-recorded content like videos on social media. This real-time generation complicates the application of age ratings, as the content cannot be easily categorized or pre-filtered. The dynamic and unpredictable nature of GAI content necessitates a more nuanced approach to content moderation and parental control. Static systems based on fixed categories or age ratings fall short in addressing the complexities of real-time generated content. Consequently, a more sophisticated and adaptable method is needed to ensure the safety and appropriateness of content for children. We propose a parent-AI collaborative system, where parents act as intermediaries between GAI and their children, evaluating, adjusting, and deciding on the output generated by GAI on sensitive topics. This collaborative system should also adapt automatically to the parents’ adjustment history, enhancing its ability to filter and restrict content autonomously over time. Through this process, parents effectively fine-tune the content filtering model in GAI, ensuring a personalized and responsive approach to content management that aligns with their children’s developmental stages and individual needs. Furthermore, given the new functionalities involved in GAI, it is imperative to involve both parents and children in co-designing the controls. A collaborative design process will help better understand the needs and dynamics from both perspectives, leading to more effective and mutually agreed-upon content management solutions.
@@ 7 Conclusion
Through Reddit content analysis and 20 interviews with teenagers and parents, we identified a major gap in parental awareness about how children use GAI—for emotional support and virtual relationships. Parents tend to worry about data privacy and misinformation, while teenagers are concerned GAI usage in social interactions, such as addiction to virtual relationships with chatbots, the distress caused by GAI-generated content in social groups, and privacy breaches involving the unauthorized use of their personal data for creating or manipulating digital content. The absence of robust parental controls further complicates this landscape, making it difficult for parents to provide effective oversight or mediation. Our study underscores the need for GAI platforms to enhance safety features and for clearer, more accessible communication tools to help parents understand and navigate the risks associated with their children’s use of these technologies. Based on our findings, we have offered design recommendations for improving parental controls on GAI platforms.

### v2

@@ 6.1 Design Implications
There are existing AI risk management frameworks, policies, and regulations for children, such as those provided by the United Nations Children’s Fund (UNICEF), the United Nations Convention on the Rights of the Child (UNCRC), and the Age Appropriate Design Code (AADC) [13]. However, finding this compliance information for specific products is difficult for users, especially children. We recommend that GAI practitioners combine these AI risk management frameworks with users’ perceptions of risks to create an easy-to-understand child risk and protection disclosure on product pages. User feedback and priorities on GAI risks and protections can guide the design of these disclosures, influencing the sequence and format of information. Additionally, this feedback can inform policymakers in creating more effective and user-centered regulations. To make this disclosure easily understandable for both parents and children, GAI practitioners could draw inspiration from label designs used for children’s products in everyday life. Emami-Naeini et al. have prototyped security and privacy labels for IoT devices, which could serve as a model for creating similar labels for GAI platforms. This would help parents and children by providing clear, accessible information about the safety and privacy measures in place for each platform, lowering the barrier for them to make informed decisions. To improve engagement, these pages could be made more interactive or visually engaging, such as through infographics or short videos that highlight key risks and protective measures.
@@
While existing parental control features are useful, they are not entirely suitable for managing appropriate content on Generative AI (GAI) platforms. The primary challenge stems from the nature of GAI, which generates real-time outputs rather than providing pre-recorded content like videos on social media. This real-time generation complicates the application of age ratings, as the content cannot be easily categorized or pre-filtered. The dynamic and unpredictable nature of GAI content necessitates a more nuanced approach to content moderation and parental control. Static systems based on fixed categories or age ratings fall short in addressing the complexities of real-time generated content. Consequently, a more sophisticated and adaptable method is needed to ensure the safety and appropriateness of content for children. We propose a parent-AI collaborative system, where parents act as intermediaries between GAI and their children, evaluating, adjusting, and deciding on the output generated by GAI on sensitive topics. The implementation of these protections can be tailored to fit the platform’s specific functionality. For platforms like Character.ai, instead of a generic risk warning, safety features could be integrated in a way that aligns with the role of chatbot and remains in character. This collaborative system should also adapt automatically to the parents’ adjustment history, enhancing its ability to filter and restrict content autonomously over time. Through this process, parents effectively fine-tune the content filtering model in GAI, ensuring a personalized and responsive approach to content management that aligns with their children’s developmental stages and individual needs. Furthermore, given the new functionalities involved in GAI, it is imperative to involve both parents and children in co-designing the controls. A collaborative design process will help better understand the needs and dynamics from both perspectives, leading to more effective and mutually agreed-upon content management solutions.
@@ 7 Conclusion
Through Reddit content analysis and 20 interviews with teenagers and parents, we found that teenagers use GAI for diverse purposes, ranging from emotional support and social interactions to engaging in risky behaviors, such as manipulating chatbots. Additionally, there is a considerable gap between how parents perceive the risks of GAI and how teenagers actually use these tools. Existing parental mediation strategies are often inadequate for managing teenagers’ real-time interactions with GAI. This highlights the need for GAI platforms to implement more comprehensive, age-appropriate content moderation systems that can effectively address these risky behaviors. Specifically, platforms should develop features to mitigate addiction risks, particularly in cases where teenagers form unhealthy dependencies on virtual relationships with AI. Moreover, GAI platforms must offer parent-focused resources that clearly outline both the capabilities and potential risks of GAI, using accessible and straightforward language. Finally, enhancing parental control systems with real-time monitoring and customizable settings will empower parents to better manage their teenagers’ engagement with GAI, ensuring a safer and more informed use of these technologies.
Appendix A Meta-Review
The following meta-review was prepared by the program committee for the 2025 IEEE Symposium on Security and Privacy (S&P) as part of the review process as detailed in the call for papers.
A.1 Summary
This paper examines the risks for children and parental control challenges related to the use of Generative AI (GAI) by performing a content analysis on Reddit and conducting semi-structured online interviews with parents and children and found varying perceptions of GAI, such as a lack of parental awareness regarding their children’s use of GAI, and differing concerns about the associated risks.
A.2 Scientific Contributions
  • Provides a Valuable Step Forward in an Established Field.
  • Identifies an Impactful Vulnerability.
A.3 Reasons for Acceptance
  1. 1.
    The paper offers a valuable step forward in an established field by providing insights into the differing perceptions of GAI between parents and children. It enhances the existing body of work on digital safety for children by identifying varying concerns about the associated risks.
  2. 2.
    The paper identifies an impactful vulnerability in the lack of parental awareness about their children’s use of GAI, which raises attention to unsupervised interactions.